# A new class of non-Einstein pp-wave solutions to quadratic gravity

### v1

@@ Abstract
We obtain a new family of exact vacuum solutions to quadratic gravity that describe pp-waves with two-dimensional wave surfaces that can have any prescribed constant curvature. When the wave surfaces are flat we recover the Peres waves obtained by Madsen, a subset of which forms precisely the vacuum pp-waves of general relativity. If, on the other hand, the wave surfaces have non-vanishing constant curvature then all our solutions are non-Einstein (i.e. they do not solve Einstein’s equations in vacuum, with or without cosmological constant) and we find that the curvature is linearly related to the value of the cosmological constant. We show that the vacuum field equations reduce to a simple linear biharmonic equation on the curved wave surfaces, and as consequence, the general solution can be written down. We also provide some simple explicit examples.
@@ I Introduction
Einstein spacetimes, i.e. vacuum solutions to general relativity with or without cosmological constant, are trivially solutions in QG. Other than these, however, not many exact solutions to QG are known, due to the complexity of the highly nonlinear fourth-order field equations. Some explicit non-trivial solutions have been obtained in [3, 4, 5, 6, 7, 8], all of which belong to the Kundt class or are conformal to it. Moreover, all of these solutions have a constant Ricci scalar, which simplifies the field equations enormously. Spherically symmetric solutions have been characterized recently in terms of infinite power series [9, 10].
In this paper we introduce a new family of non-Einstein exact pp-wave solutions to QG. We define a pp-wave in the general sense, as a Lorentzian metric that admits a covariantly constant null vector field (CCNV). In general relativity it is then a consequence of Einstein’s field equations that (i) the wave surfaces of a vacuum pp-wave are always flat in standard coordinates, and that (ii) pp-waves are not reconcilable with a non-vanishing cosmological constant [^1] We note, however, that the related Siklos solutions, which are (in some sense) conformal to pp-waves, can be interpreted as exact gravitational waves traveling on an AdS background [11]. Their generalization to QG has been studied in [4].. The solutions that we will present here, on the other hand, show that QG allows for vacuum pp-waves with wave surfaces of any constant curvature, where this curvature is proportional to the value of the cosmological constant. We show that the field equations reduce to a simple linear biharmonic equation on the wave surfaces, for which the general solution can be written down. When the wave surfaces are assumed to be flat, we recover the classical pp-waves of GR and their QG generalizations [3].
@@ II QG Action and Field Equations
as well as signature convention ($+$)(-$+$)( - + + + ). In 4 spacetime dimensions, it follows from the Chern-Gauss-Bonnet theorem and the Ricci decomposition of the curvature tensor that the $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ term and the $R_{\mu\nu}R^{\mu\nu}$ term can simultaneously be eliminated in favour of a single $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ term involving the Weyl tensor,
@@
The solutions that we will present in this paper belong to the class of pp-waves (plane-fronted waves with parallel rays). In general, pp-waves are Lorentzian metrics defined by the property that they admit a covariantly constant null vector field (CCNV). They were described for the first time by Brinkmann [12] and have been studied in detail in e.g. [13, 14] in the context of GR (see section 24.5 in [15] for a summary). For an algebraic classification independent of any field equations, see [16], and see [17] for an enlightening recent review. Roughly speaking pp-waves can be interpreted as exact gravitational waves travelling on a Minkowski background.
It is well-known that a pp-wave metric can always be expressed in suitable local coordinates as
@@ IV Vacuum pp-waves in quadratic gravity
Clearly this ansatz falls under the general definition of a pp-wave given above. However, whereas the classes of metrics $(B)$ and $(C)$ can never be vacuum solutions to Einstein’s equation, as we will confirm explicitly below, we will find that they do solve the vacuum equations of QG, provided that the function $H$ satisfies the differential equation $\Delta^{2}H=0$, where $\Delta$ is the 2D Laplace-Beltrami operator on the wave surfaces.
Metrics in class $(A)$, on the other hand, are precisely metrics of the form (16), and it is well-known that there is a subset of this class that solves Einstein’s field equations in vacuum, characterized by $\delta^{ab}\partial_{a}\partial_{b}H=0$. Nevertheless, in QG the class of solutions of this form is larger. This case was already covered in [3] and we will only review it briefly below.
Before we move on to the analysis of the field equations, it is worth pointing out that our ansatz is in some sense orthogonal to the gyratonic pp-wave ansatz recently analyzed in [7] for quadratic gravity, where the appropriate generalization of (16) is the presence of non-vanishing $W_{a}$ as in (15). Our present approach, on the other hand, is to stick to $W_{a}=0$, while instead generalizing the flat $\delta_{ab}$ in (16) to a curved $h_{ab}$ as in (15).
@@ IV.1 Flat wave surfaces
where $H$ is an arbitrary smooth function. The only non-vanishing component of the Ricci tensor is given by
@@
and the only non-vanishing component of the Bach tensor is given by
@@
where we recall that $\kappa$ is defined as in (14). This agrees [^6] Our $\kappa$ is related to $K$ in [3] by $K^{2}=-1/(2\kappa)$. with the results in [3]. The general solution to this equation has been constructed in [3] as well and there’s no need to repeat that analysis here.
@@ IV.2.1 The field equations
The non-vanishing Christoffel symbols are given by
$\Gamma^{v}_{uu}$ $=\frac{\partial H}{\partial u},$ $\Gamma^{v}_{u\theta}$ $=\Gamma^{v}_{\theta u}=\frac{\partial H}{\partial\theta},$ (25)
$\Gamma^{v}_{u\phi}$ $=\Gamma^{v}_{\phi u}=\frac{\partial H}{\partial\phi},$ $\Gamma^{\theta}_{uu}$ $=-\frac{1}{r^{2}}\frac{\partial H}{\partial\theta},$ (26)
$\Gamma^{\theta}_{\phi\phi}$ $=-\cos{\theta}\sin{\theta},$ $\Gamma^{\phi}_{uu}$ $=-\frac{1}{r^{2}\sin^{2}{\theta}}\frac{\partial H}{\partial\phi},$ (27)
$\Gamma^{\phi}_{\theta\phi}$ $=\Gamma^{\phi}_{\phi\theta}=\frac{1}{\tan{\theta}},$ (28)
and hence the non-vanishing components of the Ricci tensor are given by:
@@
$B_{uu}$ $=-\frac{1}{r^{4}}\left(\frac{1}{2}\Delta_{\mathbb{S}}^{2}+\frac{1}{3}\Delta_{\mathbb{S}}+\frac{1}{3}\right)H$
$B_{uv}$ $=B_{vu}=-\frac{1}{6r^{4}},$
@@
So let’s assume, throughout the remainder of this section, that $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ and $\Lambda$ are related as above. The only remaining nontrivial component of the field equations (13) reads
@@ IV.2.2 Globally spherical solutions to $\Delta_{\mathbb{S}}^{2}H=0$
In order to find the general solution, we return to the formulation in (45). It is a well-known result that there are no harmonic functions on the sphere, other than the constant ones, so one might wonder if the class of (smooth) biharmonic functions is any larger. This turns out not to be the case. We will prove momentarily that any biharmonic function on the sphere is necessarily constant. This assumes, however, that the wave surfaces are globally spherical. In contrast, we will show in Section IV.2.3 below that, locally, non-trivial solutions do exist. We construct the general solution and provide a simple example.
@@
The term by term differentiation is justified by virtue of the fact that $H$ is smooth (see Appendix A for a proof sketch). Since the $Y_{\ell m}$ form an orthonormal basis, this can only be satisfied if all the coefficients vanish, $H_{\ell m}\ell^{2}(\ell+1)^{2}=0$. It follows therefore that $H_{\ell m}=0$ whenever $\ell\neq 0$, and hence that $H=H_{00}Y_{00}=$ constant.
@@ IV.2.4 A simple example
Because of the branch-cut of the natural logarithm, this function cannot be extended to a smooth function on the whole sphere, but for $\theta\in(0,\pi)$, say, it is harmonic and hence in particular biharmonic, i.e. it satisfies the field equations (45). Hence, for $r$ given by (40), this example yields an exact vacuum solution to the field equations of QG with a value of the cosmological constant given by (43). Note that it is not a vacuum solution to Einstein’s equations, since the metric is not Einstein, as shown above.
@@ IV.3.1 The field equations
where $H$ is again an arbitrary smooth function. The nonvanishing Christoffel symbols of this metric are given by
$\Gamma^{v}_{uu}$ $=\frac{\partial H}{\partial u},\qquad$ $\Gamma^{v}_{ux}$ $=\Gamma^{v}_{xu}=\frac{\partial H}{\partial x},$
$\Gamma^{v}_{uy}$ $=\Gamma^{v}_{yu}=\frac{\partial H}{\partial y},\qquad$ $\Gamma^{x}_{uu}$ $=-\frac{x^{2}}{r^{2}}\frac{\partial H}{\partial x},$
$\Gamma^{x}_{xx}$ $=-\frac{1}{x},\qquad$ $\Gamma^{x}_{yy}$ $=\frac{1}{x},$
$\Gamma^{y}_{uu}$ $=-\frac{x^{2}}{r^{2}}\frac{\partial H}{\partial y},\qquad$ $\Gamma^{y}_{xy}$ $=\Gamma^{y}_{yx}=-\frac{1}{x}.$ (64)
The nonzero components of the Ricci tensor are given by:
@@
$B_{uu}$ $=-\frac{1}{r^{4}}\left[\frac{1}{2}\Delta_{\mathbb{H}}^{2}-\frac{1}{3}\Delta_{\mathbb{H}}+\frac{1}{3}\right]H$ (71)
$B_{uv}$ $=B_{vu}=-\frac{1}{6r^{4}},$
$B_{xx}$ $=B_{yy}=\frac{1}{6r^{2}x^{2}},$
@@
So let’s assume, throughout the remainder of this section, that $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ and $\Lambda$ are related as above. The only remaining nontrivial component of the field equations (13) reads
@@
Formally, this is exactly the same equation (45) as the one we obtained in the case of spherical wave surfaces. The solutions are again precisely the biharmonic functions, in this case on hyperbolic space. Again we note that any harmonic function, i.e. $\Delta_{\mathbb{H}}H=0$, is automatically biharmonic, which, in constrast to the spherical case, already provides a large range of solutions that are defined on the entire hyperbolic plane. The class of biharmonic functions on hyperbolic space is strictly larger, though, as we discuss below.
@@ IV.3.2 The general solution to $\Delta_{\mathbb{H}}^{2}H=0$
Note that the second integral (91) always exists, since the factor multiplying $h(\phi_{0})$ in the integrand is $L^{1}$, while $h$ is $L^{\infty}$, and the same is true for the integral (87) that defines $\phi$. For the first integral (90), however, the $1/(1-\rho^{2})^{2}$ behaviour of $1/f$ complicates things. A sufficient condition for the integral to exist is that $\rho\phi(\rho,\phi)/f(\rho,\phi)$ is bounded on the unit disk, but we haven’t been able to find a simple condition on $\phi$ alone that enforces such behaviour. In fact, we expect non-vanishing functions $\phi$ satisfying said sufficient condition to be rare, which would indicate that most (but certainly not all, as shown below) solutions to $\Delta_{\mathbb{H}}^{2}H=0$ are in fact harmonic, $\Delta_{\mathbb{H}}H=0$. Such solutions can be obtained from the general solution above simply by setting $\phi=0$ (equivalently, $\Phi=0$).
@@ V Discussion
First, (i) if the cosmological constant is positive then the wave surfaces must be spherical. In this case we have shown that the solutions should be understood only locally; indeed if the wave surfaces are assumed to be globally spherical then only a single rather trivial solution exists, by virtue of the fact that there are no non-constant smooth biharmonic functions on the sphere.
@@
As mentioned before, our approach has in some sense been complementary to the approach in [7], in which non-vanishing gyratonic terms in the pp-wave metric were considered, while leaving the wave surfaces flat—in contrast to the approach we have taken here, generalizing the wave surfaces by allowing them to be curved, while assuming instead that the gyratonic terms all vanish. In future work, it would be interesting to investigate the scenario in which both gyratonic terms and curved wave surfaces are present. Similarly, it would be very much of interest to see whether the results derived in the present work can be generalized to wave surfaces of arbitrary—not necessarily constant—curvature.

### v2

# New class of non-Einstein pp-wave solutions to quadratic gravity

@@ Abstract
We obtain a new family of exact vacuum solutions to quadratic gravity that describe pp-waves with two-dimensional wave surfaces that can have any prescribed constant curvature. When the wave surfaces are flat we recover the Peres waves obtained by Madsen, a subset of which forms precisely the vacuum pp-waves of general relativity. If, on the other hand, the wave surfaces have nonvanishing constant curvature then all of our solutions are non-Einstein (i.e. they do not solve Einstein’s equations in vacuum, with or without cosmological constant) and we find that the curvature is linearly related to the value of the cosmological constant. We show that the vacuum field equations reduce to a simple linear biharmonic equation on the curved wave surfaces, and as consequence, the general solution can be written down. We also provide some simple explicit examples.
@@ I Introduction
Einstein spacetimes, i.e. vacuum solutions to general relativity with or without cosmological constant, are trivially solutions in QG. Other than these, however, not many exact solutions to QG are known, due to the complexity of the highly nonlinear fourth-order field equations. Some explicit nontrivial solutions have been obtained in [3, 4, 5, 6, 7, 8], all of which belong to the Kundt class or are conformal to it. Moreover, all of these solutions have a constant Ricci scalar, which simplifies the field equations enormously. Spherically symmetric solutions have been investigated numerically and by using perturbative methods and series expansions [9, 10, 11, 12, 13].
In this paper we introduce a new family of non-Einstein exact pp-wave solutions to QG. We define a pp-wave in the general sense, as a Lorentzian metric that admits a covariantly constant null vector field (CCNV). In general relativity it is then a consequence of Einstein’s field equations that (i) the wave surfaces of a vacuum pp-wave are always flat in standard coordinates, and that (ii) pp-waves are not reconcilable with a nonvanishing cosmological constant [^1] We note, however, that the related Siklos solutions, which are (in some sense) conformal to pp-waves, can be interpreted as exact gravitational waves traveling on an AdS background [14]. Their generalization to QG has been studied in [4].. The solutions that we will present here, on the other hand, show that QG allows for vacuum pp-waves with wave surfaces of any constant curvature, where this curvature is proportional to the value of the cosmological constant. We show that the field equations reduce to a simple linear biharmonic equation on the wave surfaces, for which the general solution can be written down. When the wave surfaces are assumed to be flat, we recover the classical pp-waves of GR and their QG generalizations [3].
@@ II QG Action and Field Equations
as well as signature convention ($+$)(-$+$)( - + + + ). In four spacetime dimensions, it follows from the Chern-Gauss-Bonnet theorem and the Ricci decomposition of the curvature tensor that the $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ term and the $R_{\mu\nu}R^{\mu\nu}$ term can simultaneously be eliminated in favor of a single $C_{\mu\nu\rho\sigma}C^{\mu\nu\rho\sigma}$ term involving the Weyl tensor,
@@
The solutions that we will present in this paper belong to the class of pp-waves (plane-fronted waves with parallel rays). In general, pp-waves are Lorentzian metrics defined by the property that they admit a CCNV. They were described for the first time by Brinkmann [15] and have been studied in detail in e.g. [16, 17] in the context of GR (see section 24.5 in [18] for a summary). For an algebraic classification independent of any field equations, see [19], and see [20] for an enlightening recent review. Roughly speaking pp-waves can be interpreted as exact gravitational waves traveling on a Minkowski background.
It is well known that a pp-wave metric can always be expressed in suitable local coordinates as
@@ IV Vacuum pp-waves in quadratic gravity
Clearly this ansatz falls under the general definition of a pp-wave given above. However, whereas the classes of metrics $(B)$ and $(C)$ can never be vacuum solutions to Einstein’s equations, as we will confirm explicitly below, we will find that they do solve the vacuum equations of QG, provided that the function $H$ satisfies the differential equation $\Delta^{2}H=0$, where $\Delta$ is the 2D Laplace-Beltrami operator on the wave surfaces. Interestingly, related GR solutions with nonvanishing cosmological constant do exist. These represent impulsive gravitational waves generated by null point sources, and have wave surfaces of constant curvature [22, 23].
Metrics in class $(A)$, on the other hand, are precisely metrics of the form (16), and it is well known that there is a subset of this class that solves Einstein’s field equations in vacuum, characterized by $\delta^{ab}\partial_{a}\partial_{b}H=0$. Nevertheless, in QG the class of solutions of this form is larger. This case was already covered in [3] and we will only review it briefly below.
Before we move on to the analysis of the field equations, it is worth pointing out that our ansatz is in some sense orthogonal to the gyratonic pp-wave ansatz recently analyzed in [7] for quadratic gravity, where the appropriate generalization of (16) is the presence of nonvanishing $W_{a}$ as in (15). Our present approach, on the other hand, is to stick to $W_{a}=0$, while instead generalizing the flat $\delta_{ab}$ in (16) to a curved $h_{ab}$ as in (15).
@@ IV.1 Flat wave surfaces
where $H$ is an arbitrary smooth function. The only nonvanishing component of the Ricci tensor is given by
@@
and the only nonvanishing component of the Bach tensor is given by
@@
where we recall that $\kappa$ is defined as in (14). This agrees [^6] Our $\kappa$ is related to $K$ in [3] by $K^{2}=-1/(2\kappa)$. with the results in [3]. The general solution to this equation has been constructed in [3] as well and there is no need to repeat that analysis here.
@@ IV.2.1 The field equations
The nonvanishing components of the Ricci tensor are then given by:
@@
$B_{uu}$ $=-\frac{1}{r^{4}}\left(\frac{1}{2}\Delta_{\mathbb{S}}^{2}+\frac{1}{3}\Delta_{\mathbb{S}}+\frac{1}{3}\right)H$ (32)
$B_{uv}$ $=B_{vu}=-\frac{1}{6r^{4}},$ (33)
@@
So let us assume, throughout the remainder of this section, that $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ and $\Lambda$ are related as above. The only remaining nontrivial component of the field equations (13) reads
@@ IV.2.2 Globally spherical solutions to $\Delta_{\mathbb{S}}^{2}H=0$
In order to find the general solution, we return to the formulation in (43). It is a well-known result that there are no harmonic functions on the sphere, other than the constant ones, so one might wonder if the class of (smooth) biharmonic functions is any larger. This turns out not to be the case. We will prove momentarily that any biharmonic function on the sphere is necessarily constant. This assumes, however, that the wave surfaces are globally spherical. In contrast, we will show in Section IV.2.3 below that, locally, nontrivial solutions do exist. We construct the general solution and provide a simple example.
@@
The term-by-term differentiation is justified by virtue of the fact that $H$ is smooth (see Appendix A for a proof sketch). Since the $Y_{\ell m}$ form an orthonormal basis, this can only be satisfied if all the coefficients vanish, $H_{\ell m}\ell^{2}(\ell+1)^{2}=0$. It follows therefore that $H_{\ell m}=0$ whenever $\ell\neq 0$, and hence that $H=H_{00}Y_{00}=$ constant.
@@ IV.2.4 A simple example
Because of the branch cut of the natural logarithm, this function cannot be extended to a smooth function on the whole sphere, but for $\theta\in(0,\pi)$, say, it is harmonic and hence in particular biharmonic, i.e. it satisfies the field equations (43). Hence, for $r$ given by (38), this example yields an exact vacuum solution to the field equations of QG with a value of the cosmological constant given by (41). Note that it is not a vacuum solution to Einstein’s equations, since the metric is not Einstein, as shown above.
@@ IV.3.1 The field equations
where $H$ is again an arbitrary smooth function. The nonzero components of the Ricci tensor are then given by:
@@
$B_{uu}$ $=-\frac{1}{r^{4}}\left[\frac{1}{2}\Delta_{\mathbb{H}}^{2}-\frac{1}{3}\Delta_{\mathbb{H}}+\frac{1}{3}\right]H,$ (68)
$B_{uv}$ $=B_{vu}=-\frac{1}{6r^{4}},$ (69)
$B_{xx}$ $=B_{yy}=\frac{1}{6r^{2}x^{2}},$ (70)
@@
So let us assume, throughout the remainder of this section, that $\alpha^{\prime},\beta^{\prime},\gamma^{\prime}$ and $\Lambda$ are related as above. The only remaining nontrivial component of the field equations (13) reads
@@
Formally, this is exactly the same equation (43) as the one we obtained in the case of spherical wave surfaces. The solutions are again precisely the biharmonic functions, in this case on hyperbolic space. Again we note that any harmonic function, i.e. $\Delta_{\mathbb{H}}H=0$, is automatically biharmonic, which, in contrast to the spherical case, already provides a large range of solutions that are defined on the entire hyperbolic plane. The class of biharmonic functions on hyperbolic space is strictly larger, though, as we discuss below.
@@ IV.3.2 The general solution to $\Delta_{\mathbb{H}}^{2}H=0$
Note that the second integral (88) always exists, since the factor multiplying $h(\phi_{0})$ in the integrand is $L^{1}$, while $h$ is $L^{\infty}$, and the same is true for the integral (85) that defines $\phi$. For the first integral (87), however, the $1/(1-\rho^{2})^{2}$ behavior of $1/f$ complicates things. A sufficient condition for the integral to exist is that $\rho\phi(\rho,\phi)/f(\rho,\phi)$ is bounded on the unit disk, but we have not been able to find a simple condition on $\phi$ alone that enforces such behavior. In fact, we expect nonvanishing functions $\phi$ satisfying said sufficient condition to be rare, which would indicate that most (but certainly not all, as shown below) solutions to $\Delta_{\mathbb{H}}^{2}H=0$ are in fact harmonic, $\Delta_{\mathbb{H}}H=0$. Such solutions can be obtained from the general solution above simply by setting $\phi=0$ (equivalently, $\Phi=0$).
@@ IV.3.3 A simple example
We end this section with a final remark about the reduction of the field equations to
$\left(2\kappa\Delta_{\mathbb{E}}^{2}-\Delta_{\mathbb{E}}\right)H=0,\quad\Delta_{\mathbb{S}}^{2}H=0,\quad\Delta_{\mathbb{H}}^{2}H=0,$ (91)
in the case of flat, spherical, and hyperbolic wave surfaces, respectively. We have discussed this on a case-by-case basis. Alternatively, the result for all three cases can be proven simultaneously by choosing coordinates on the wave surfaces corresponding to the metric
$\text{d}s^{2}=\frac{d\rho^{2}}{1-k\rho^{2}}+\rho^{2}d\theta^{2}.$ (92)
This covers the flat case ($k=0$), the spherical case $(k=1)$, and the hyperbolic case ($k=-1)$.
@@ V Discussion
First, (i) if the cosmological constant is positive then the wave surfaces must be spherical. In this case we have shown that the solutions should be understood only locally; indeed if the wave surfaces are assumed to be globally spherical then only a single rather trivial solution exists, by virtue of the fact that there are no nonconstant smooth biharmonic functions on the sphere.
@@
As mentioned before, our approach has in some sense been complementary to the approach in [7], in which nonvanishing gyratonic terms in the pp-wave metric were considered, while leaving the wave surfaces flat—in contrast to the approach we have taken here, generalizing the wave surfaces by allowing them to be curved, while assuming instead that the gyratonic terms all vanish. In future work, it would be interesting to investigate the scenario in which both gyratonic terms and curved wave surfaces are present. Similarly, it would be very much of interest to see whether the results derived in the present work can be generalized to wave surfaces of arbitrary—not necessarily constant—curvature.